OpenStudy (anonymous):

If f(x) = x2 + 2, find f(4).

OpenStudy (anonymous):

is this f(x) = x^2+2 or f(x) = 2x + 2 ? Either way you, replace x with 4 when it says If f(x) =..., find f(4) f(x) = x^2 + 2 4^2 +2 16 + 2 = 18 f(x) = 2x + 2 2(4) + 2 8 + 2 = 10

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

as long as the x and two stay on the same side with each other it doesn't matter which order you place it. hope that wasn't confusing.